OpenStudy (anonymous):

The cost of four scarves and six hats is $52.00. The cost of two hats is $1.00 more than the cost of one scarf. a. What is a system of equations to represent the solution? b. What is the cost of one scarf? c. What is the cost of one hat?

OpenStudy (anonymous):

Let's say that the cost of a scarf= s and the cost of a hat = h. a. 4s + 6h = 52 ; 2h = s + 1 so, solve for s using the second equation. s = 2h - 1 Then substitute into the first equation. 4(2h - 1) + 6h = 52 Solve for h. 8h - 4 + 6h = 52 14h - 4 = 52 14h = 56 h = 56/14 h = 4 Then solve for s by substituting the h value (14) into s = 2h -1 s = 2(4) - 1 s = 7

OpenStudy (anonymous):

Is this the answer?